OpenStudy (anonymous):

Find f(-1) and f(2) for the following function. f(x)= 3/x^2-4

OpenStudy (anonymous):

\[f(\clubsuit)=\frac{3}{\clubsuit^2-4}\] \[f(-1)=\frac{3}{(-1)^2-4}\] \[f(2)=\frac{3}{2^2-4}\]

OpenStudy (anonymous):

second one is undefined since you cannot divide by zero

OpenStudy (anonymous):

so is that the infinity symbol?